In a neighborhood in a large city, the number of kids in a family can be modeled by the following box model: [0,0,1,1,1,1,1,2,2,2]. Use this box model to answer questions 8 to 11.
Calculate the average of this box model.

Part A options:

A)

1.8

B)

2.1

C)

1.1

D)

0.7

E)

0.5

If you were to select 50 families from this neighborhood, how many kids do you expect to find?

Part B options:

A)

50

B)

25

C)

68.75

D)

35

E)

55

Suppose the standard deviation of the box model is 0.70. What is the standard error of the number of kids from this sample of 50 families?

Part C options:

A)

7.7782

B)

4.9497

C)

35

D)

5.1913

E)

3.2404

What is the probability that the total number of kids in the sample is below 53?

Part D options:

A)

-0.4

B)

0.6554

C)

0.61

D)

0.3446

E)

0.7291
Answer #1

The box model is [0,0,1,1,1,1,1,2,2,2]

i.e,

X f(X)=Frequency
0 2
1 5
2 3
Total (n) 10

(8) Thus the average of this box model is :

= (0*2+1*5+3*2)/(2+5+3)= 1.1 (i.e, option C)

(9) For a 50 family of neighborhood The expected no. of children would be 50*E(Z), Where E(Z) is the expected value of children in a family, which is equal to sample mean, i.e, 1.1, Thus 50*E(Z)= 55 (i.e, option E)

(10) The SD is given to be 0.7, The sum of the no. of children in the set of 50 family is given as SD* (this is the sd of sums of all 50 families)= * SD = 7.07 * 0.7 = 4.9497(i.e, option B)

(11) We are to find the probability that P(Z<53)= P(Z-55/(SD*sqrt(n)) < 53-55/(SD*sqrt(n))) = (Normalization is aproximated as the no. of families(=50) is quite high)= 0.3446 (i.e, option D)